# LetsGo: Large-Scale Garage Modeling and Rendering via LiDAR-Assisted Gaussian Primitives

### v2

@@ 5.3 Ablation Study
Figure 11: Qualitative ablation study on the random-resolution-level (RRL) training strategy reveals that our RRL approach effectively addresses the overfitting issue at a certain level, influenced by the camera distribution.
6 GarageWorld
Figure 12: Autonomous vehicle parking. Our diverse garage scenes facilitate training algorithms for generating parking trajectories under different scenarios. When guiding the vehicle to the parking space, our garage model allows real-time and wide-FOV rendering of the environment, capturing drivable area and obstables, thereby enhancing safe parking capabilities.
We introduce the first large-scale garage dataset, GarageWorld, captured with our Polar device. This dataset comprises eight large garages, featuring six underground garages, one outdoor parking lot, and one multi-floor indoor garage. These environments present diverse and intricate geometric structures, including sloped surfaces with distorted lines, internal circular paths, spiral ramps, and parked vehicles, as detailed in Table 1. Figure 5 visualizes the 3D models and rendered images of a subsection of these garages using our method. Our rendering pipeline and the GarageWorld dataset are designed to support a wide range of applications, which are further detailed in the following sections.
@@ 6 GarageWorld
Our developed 3D reconstruction of large-scale garage scenes with enhanced real-time rendering can support various autonomous driving algorithms. The GarageWorld dataset provides diverse driving scenarios, including tight parking spaces and complex geometric layouts, which are instrumental for training algorithms in autonomous parking and navigation path planning. Equipped with sensors such as LiDAR, IMU, and cameras, autonomous vehicles utilize these 3D structures for effective sensor fusion, enhancing their operational efficiency in varied garage environments. Additionally, our LetsGo pipeline further aids vehicle parking by supplementing sensor data, particularly in low-light garage conditions where camera and radar readings may be compromised. Our LiDAR-assisted Gaussian primitives allow rigorous testing in a virtual setting, prompting safe and efficient parking solutions. Fig. 12 illustrates how our approach generates rear-view images, allowing vehicles to accurately estimate distances and navigate tight spaces, thereby avoiding collisions and accelerating the development and deployment of safe autonomous driving technologies.
@@
Our LiDAR-assisted garage modeling and rendering significantly enhance autonomous vehicle localization and navigation by providing accurate 3D references that complement real-time sensor readings. In challenging indoor or underground garages, where GPS signals are typically unavailable, traditional localization methods encounter difficulties. Confined spaces and structured obstacles can cause interference for LiDAR and radar sensors, and multipath reflections can distort sensor readings. Low lighting conditions also reduce camera visibility, complicating vision-based localization. By integrating our Polar scanner with LiDAR-assisted Gaussian primitives, we offer accurately colored 3D references that enhance feature recognition and matching, significantly improving localization and navigation reliability. Figure 13 presents an example of real-time localization and navigation. Our system correlates captured images with our colored 3D models to compute relative poses and generate optimal paths through complex environments. During driving, our approach augments camera views and extends the vehicle’s field of view, ensuring comprehensive environmental awareness and safer navigation. Additionally, our web-based rendering engine represents expansive 3D garage maps and navigation trajectories on lightweight devices with limited computational capacity, delivering smooth real-time interactions and superior rendering performance.
Figure 13: Real-time localization & navigation in challenging garage environments. Our colored 3D model facilitates precise vehicle camera localization and optimal path navigation, particularly in low-light garage conditions, ensuring safe driving. The lightweight web-based rendering ensures deployability in vehicles with limited computing resources.
@@
Our reconstructed garage datasets and real-time rendering capabilities significantly enhance VFX production by providing a robust foundation for creating realistic backgrounds and seamlessly integrating CGI elements into live-action footage. Inspired by high-adrenaline sequences like the highway chase in ”The Matrix Reloaded,” our LetsGo pipeline excels in modeling and rendering complex garage scenes. This capability allows for precise control over cinematic elements such as exposure time and motion blur, which are often challenging to manage on set. We created a VFX video production to showcase our technology using our reconstructed garage models (Fig. 14) in the supplementary video. We downloaded a video depicting highway chase scenes with dynamic viewing angles of a rapidly moving motorcycle. By extracting and aligning camera trajectories with our models and synthesizing motion blur, we created immersive visual effects that convincingly depict a chase within a garage environment. Our approach supports the exploration of diverse camera angles and compositions in real-time while ensuring visual consistency between real and virtual components, significantly reducing the need for reshoots and streamlining the production process. This integration empowers VFX artists to enhance the believability of scenes while conserving time and resources.
Figure 14: VFX demonstration. Through an analysis of the animation in the reference video, we extract the poses of several keyframes, which enable our system’s renderer to generate corresponding video segments. Our 3D garage modeling and rendering also enables motion blur rendering, producing realistic visual effects.
7 Limitations and Discussions
Although our rendering results exhibit commendable realism and we showcase various applications, our workflow still possesses certain limitations. Here we present a detailed analysis and explore further potential applications.
@@ 7 Limitations and Discussions
As a method based on image rendering, our approach achieves highly realistic rendering effects, almost indistinguishable from real scenes. However, the existing pipeline does not support modifying lighting conditions. This requires us to carefully design the shooting process according to the lighting conditions, limiting the applicability of our method in various scenarios. Enabling rendering and editing of large-scale scenes under different lighting conditions is a meaningful direction that deserves further investigation. Our open-source dataset provides a foundation for the community to conduct research in these directions, allowing for advancements in this field.
@@
Furthermore, there are additional directions and applications worth exploring. Inspired by SMERF [17], one potential research direction is to investigate streaming transmission methods for Gaussian kernels, enabling the distribution and on-the-fly rendering of large-scale scene data. Additionally, there are city generation methods such as InfiniCity [45] and CityDreamer [85], which leverage Generative Adversarial Networks (GANs) [26] to achieve rapid modeling of large-scale scenes. We intend to study large-scale scene generation based on 3D Gaussian representations. This fully explicit representation can be easily integrated into existing computer graphics workflows and achieve superior rendering effects.
8 Conclusion
@@
For many of us, our daily lives begin with a safe departure from a garage and end with a safe arrival. The garage serves as the origin of our journey to innovation. This paper has contributed a handheld Polar device for data collection, a GarageWorld dataset, LiDAR-assisted Gaussian splatting for scene representation, and an LOD-based rendering technique that allows web-based rendering on consumer-level devices. Benefiting from these innovations, we successfully reconstruct various garages with diverse and challenging environments, allowing real-time rendering from any viewpoint on lightweight devices. Experimental results on the collected and two public datasets have demonstrated the effectiveness of our approach. Our GarageWorld, along with the reconstructed 3D model and real-time rendering, enables a set of applications, including training data generation and testbed for autonomous driving algorithms, real-time assistance for autonomous vehicle localization, navigation, and parking, as well as VFX production. Our current contributions mainly focus on the perception of the world, and it enables downstream recognition tasks. In the future, we will also explore garage generation, continuing to push the boundary of garage modeling and accomplishing a closure from perception, recognition, and generation.
@@ Appendix A Specifications and Calibration of Polar Device
For the intrinsic parameter calibration of the fisheye camera and the IMU sensor, we leverage the program provided by OpenCV and the Allan Variance ROS toolbox [59, 22, 21, 50, 56], respectively. The extrinsic calibration between the fisheye camera and the IMU uses the Kalibr calibration program [59]. Due to the sparsity, noisiness, and uncolored nature of the point cloud data collected by our LiDAR sensor, it is hard to establish correspondences between the LiDAR data and the images captured by the fisheye camera, making relative pose estimation between the sensors difficult. To address this issue, we introduce an additional sensor, a FARO laser scanner. This scanner provides dense, accurate, and colored point clouds and thus functions as a bridge for the relative pose estimation between the LiDAR sensor and the fisheye camera. Specifically, we adopt a checkerboard as our calibration scene and capture data from it using the FARO laser scanner and Polar device. The relative pose between the FARO laser scanner and the LiDAR sensor in our Polar device is achieved by point cloud registration, and that between the FARO laser scanner and the fisheye camera is obtained by establishing color correspondences between the point cloud captured by the FARO sensor and the fisheye image. By connecting the transformation from FARO to LiDAR and from FARO to the fisheye camera, we can finally calculate the extrinsic parameters between the LiDAR and the fisheye camera. We repeat this process twelve times in our implementation to reduce experimental error. It is important to note that FARO was only used to calibrate between the fisheye camera and the LiDAR before we start scanning a garage, and the process only needs to be conducted once.
@@ Appendix B Data Preprocessing and Mesh cleanup
To ensure data privacy, we anonymized vehicle license plate information recorded in the underground garage data collection. Specifically, we used image blurring to hide license plate numbers from identification. This step is essential to protect individual privacy and comply with data protection regulations. To remove the dynamic objects in the collected data, we apply Segment Anything [37] to the images. For the point cloud data, we estimate whether the points scanned at one LiDAR frame are also observed at other frames, similar to Schauer and Nüchter [63]. For mesh cleaning, we construct a kd-tree from the point cloud of LIV-SLAM and calculate the nearest neighbour distance of each mesh face (k = 1). If the distance exceeds 0.1m, we remove the face.
@@ Appendix C Details of depth calculation
To accurately calculate depth, we treat each Gaussian as a probability density function. Following [100], we get the depth of each Gaussian primitive, $d_{i}$, by calculating the expected value of depth under this probability density as:
@@
where $g_{i}$ is the Gaussian function, $\boldsymbol{x}=[x_{0},x_{1},x_{2}]^{T}$ is the coordinate of the space in the ray space, where $x_{0},x_{1}$ are the pixel coordinates, $x_{2}$ is oriented in the same direction as the pixel ray. As in [100], $t=x_{2}/l$ is the depth of $\boldsymbol{x}$ in the camera coordinate system and $l=\sqrt{x_{0}^{2}+x_{1}^{2}+1}$. We denote the center of the Gaussian in ray space as $\boldsymbol{p}=[p_{0},p_{1},p_{2}]$ and the covariance matrix as $\boldsymbol{\boldsymbol{\Sigma}}$, by defining $\boldsymbol{y}=\boldsymbol{x}-\boldsymbol{p}$, the Eqn. 9 can be simplified as:
@@
$g(\boldsymbol{y}+\boldsymbol{p})=e^{-Ay_{2}^{2}+By_{2}+C}=e^{C}\cdot e^{-Ay_{2}^{2}+By_{2}},$ (11)
@@
$\begin{split}A&=\frac{1}{2}\left(\boldsymbol{\Sigma}^{-1}\right)_{2,2}\\ B&=-\left(\boldsymbol{\Sigma}^{-1}\right)_{2,0}y_{0}-\left(\boldsymbol{\Sigma}^{-1}\right)_{2,1}y_{1}\\ C&=-\frac{1}{2}\left(\boldsymbol{\Sigma}^{-1}\right)_{0,0}y_{0}^{2}-\frac{1}{2}\left(\boldsymbol{\Sigma}^{-1}\right)_{1,1}y_{1}^{2}-\left(\boldsymbol{\Sigma}^{-1}\right)_{1,0}y_{0}y_{1}\\ .\end{split}$ (12)
@@
$R(D)=\begin{cases}\frac{1}{2\beta}D&(D<\beta)\\ 1-\frac{\beta}{2D}&(D\geq\beta)\end{cases},$ (16)

### v3

@@ 5.3. Ablation Study
Table 3. Quantitative ablation study on different input point cloud qualities and corresponding training times across two datasets. Point clouds sampled from the mesh at varying intervals and original LiDAR point clouds downsampled to 4 cm intervals. The training times are based on tests performed with an NVIDIA RTX A6000 GPU.
Point Cloud Density Arts Center Shopping Mall 3
[…]
2 cm (Mesh) 25.87 0.862 0.317 17698464 10:55:07 25.01 0.836 0.320 36042237 31:01:11
6 cm (Mesh) 25.52 0.854 0.340 2300174 02:47:02 24.73 0.827 0.340 4559012 06:35:14
8 cm (Mesh) 25.29 0.848 0.348 1306965 02:15:16 24.49 0.821 0.348 2566219 04:53:08
10 cm (Mesh) 25.08 0.843 0.356 836921 02:02:50 24.21 0.814 0.357 1627634 04:06:29
4 cm (LiDAR) 23.11 0.816 0.370 10024717 03:03:03 21.52 0.764 0.399 21162613 04:39:05
4 cm (Mesh) 25.70 0.858 0.330 4997081 02:23:27 24.91 0.832 0.331 10015908 10:58:39
@@
Input Point Cloud Qualities.
Relying solely on high-density point clouds can lead to computational inefficiencies without significantly enhancing reconstruction quality. To investigate this, we conducted an ablation study assessing the impact of varying point cloud qualities on our reconstruction results. As shown in Fig. 11 and Tab. 3, we evaluated the performance of our LOD-LiDAR-GS method using point clouds sampled from the mesh at various intervals, alongside original LiDAR point clouds downsampled to 4 cm intervals. Our findings indicate that while higher densities yield more detailed reconstructions, the benefits plateau beyond a certain threshold. Conversely, lower densities still deliver robust performance with reduced computational costs. To balance training time and reconstruction quality, we selected 4 cm intervals for sampling from the mesh. Notably, using original LiDAR point clouds directly results in poorer outcomes compared to mesh-sampled point clouds, particularly in garage environments where objects are smaller and geometrically simpler. Thus, we opted for point clouds sampled from the mesh at 4 cm intervals, which offer superior reconstruction quality while maintaining efficiency.
(a) Autonomous vehicle parking
(b) Real-time localization & navigation
(c) VFX demonstration
Figure 13. Applications of our system: (a) Autonomous vehicle parking. Our diverse garage scenes facilitate training algorithms for generating parking trajectories under different scenarios; (b) Real-time localization & navigation in challenging garage environments; (c) VFX demonstration. Through an analysis of the animation in the reference video, we extract the poses of several keyframes, which enable our system’s renderer to generate corresponding video segments and produce realistic visual effects.
@@
6. GarageWorld
We introduce the first large-scale garage dataset, GarageWorld, captured with our Polar device. This dataset comprises eight large garages, featuring six underground garages, one outdoor parking lot, and one multi-floor indoor garage. Fig. 4 visualizes the 3D models and rendered images of a subsection of these garages using our method. Our rendering pipeline and the GarageWorld dataset are designed to support a wide range of applications, which are further detailed in the following sections.
@@ 6. GarageWorld
Our developed 3D reconstruction of large-scale garage scenes with enhanced real-time rendering can support various autonomous driving algorithms. The GarageWorld dataset provides diverse driving scenarios, including tight parking spaces and complex geometric layouts, which are instrumental for training algorithms in autonomous parking and navigation path planning. Additionally, our LetsGo pipeline further aids vehicle parking by supplementing sensor data, particularly in low-light garage conditions. Our LiDAR-assisted Gaussian primitives allow rigorous testing in a virtual setting, prompting safe and efficient parking solutions. Fig. 13(a) illustrates how our approach generates rear-view images, allowing vehicles to accurately estimate distances and navigate tight spaces, thereby avoiding collisions and accelerating the development and deployment of safe autonomous driving technologies.
@@
Our LiDAR-assisted garage modeling and rendering provide accurate 3D references, significantly enhancing autonomous vehicle localization and navigation, particularly in challenging indoor or underground garages where traditional methods struggle. Fig. 13(b) presents an example of real-time localization and navigation. Additionally, our web-based rendering engine represents expansive 3D garage maps and navigation trajectories on lightweight devices with limited computational capacity, delivering smooth real-time interactions and superior rendering performance.
@@
Our reconstructed garage datasets and real-time rendering capabilities significantly enhance VFX production by providing a robust foundation for creating realistic backgrounds and seamlessly integrating CGI elements into live-action footage. Inspired by the highway chase in ”The Matrix Reloaded”, our LetsGo pipeline excels in rendering complex garage scenes, offering precise control over elements like exposure time and motion blur. We created a VFX video production to showcase our technology using our reconstructed garage models (Fig. 13(c)) in the supplementary video. We downloaded a video depicting highway chase scenes with dynamic viewing angles of a rapidly moving motorcycle. By extracting and aligning camera trajectories with our models and synthesizing motion blur, we created immersive visual effects that convincingly depict a chase within a garage environment. Our approach supports the exploration of diverse camera angles and compositions in real-time while ensuring visual consistency between real and virtual components, significantly reducing the need for reshoots and streamlining the production process.
7. Limitations and Discussions
Although our rendering results exhibit realism and we showcase various applications, our workflow still possesses certain limitations. Here we present a detailed analysis and explore further potential applications.
@@ 7. Limitations and Discussions
As a method based on image rendering, our approach achieves highly realistic rendering effects, almost indistinguishable from real scenes. However, the existing pipeline does not support modifying lighting conditions. This requires us to carefully design the shooting process according to the lighting conditions, limiting the applicability of our method in various scenarios. Enabling rendering and editing of large-scale scenes under different lighting conditions is a meaningful direction that deserves further investigation. Our open-source dataset provides a foundation for the community to conduct research in these directions, allowing for advancements in this field. Also there are switching artifacts primarily caused by loading different Gaussian levels. These artifacts can be mitigated by interpolating between adjacent Gaussian levels during switching. (Kerbl et al., 2024) successfully achieves smooth level transitions using a linear interpolation scheme. We will explore this approach and decrease the switching artifacts in future work.
@@
Furthermore, there are additional directions and applications worth exploring. Inspired by SMERF (Duckworth et al., 2024), one potential research direction is to investigate streaming transmission methods for Gaussian kernels, enabling the distribution and on-the-fly rendering of large-scale scene data. Additionally, there are city generation methods such as InfiniCity (Lin et al., 2023) and CityDreamer (Xie et al., 2024), which leverage Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) to achieve rapid modeling of large-scale scenes. We intend to study large-scale scene generation based on 3D Gaussian representations. This fully explicit representation can be easily integrated into existing computer graphics workflows and achieve superior rendering effects.
8. Conclusion
@@
For many of us, our daily lives begin with a safe departure from a garage and end with a safe arrival. The garage serves as the origin of our journey to innovation. This paper has contributed a handheld Polar device for data collection, a GarageWorld dataset, LiDAR-assisted Gaussian splatting for scene representation, and an LOD-based rendering technique that allows web-based rendering on consumer-level devices. Benefiting from these innovations, we successfully reconstruct various garages with diverse and challenging environments, allowing real-time rendering from any viewpoint on lightweight devices. Experimental results on the collected and two public datasets have demonstrated the effectiveness of our approach. Our GarageWorld, along with the reconstructed 3D model and real-time rendering, enables a set of applications, including training data generation and testbed for autonomous driving algorithms, real-time assistance for autonomous vehicle localization, navigation, and parking, as well as VFX production. Our current contributions mainly focus on the perception of the world, and enables downstream recognition tasks. In the future, we will also explore garage generation, continuing to push the boundary of garage modeling and accomplishing a closure from perception, recognition, and generation.
Acknowledgements.
The authors would like to thank Shanghai Municipal Big Data Center for coordinating the collection of garage data. This work was supported by National Key R&\&&D Program of China (2022YFF0902301), NSFC programs (61976138, 61977047), STCSM (2015F0203-000-06), and SHMEC (2019-01-07-00-01-E00003). We also acknowledge support from Shanghai Frontiers Science Center of Human-centered Artificial Intelligence (ShangHAI) and MoE Key Lab of Intelligent Perception and Human-Machine Collaboration (ShanghaiTech University).
@@ Appendix A Specifications and Calibration of Polar Device
For the intrinsic parameter calibration of the fisheye camera and the IMU sensor, we leverage the program provided by OpenCV and the Allan Variance ROS toolbox (Rehder et al., 2016; Furgale et al., 2013, 2012; Maye et al., 2013; Oth et al., 2013), respectively. The extrinsic calibration between the fisheye camera and the IMU uses the Kalibr calibration program (Rehder et al., 2016). Due to the sparsity, noisiness, and uncolored nature of the point cloud data collected by our LiDAR sensor, it is hard to establish correspondences between the LiDAR data and the images captured by the fisheye camera, making relative pose estimation between the sensors difficult. To address this issue, we introduce an additional sensor, a FARO laser scanner. This scanner provides dense, accurate, and colored point clouds and thus functions as a bridge for the relative pose estimation between the LiDAR sensor and the fisheye camera. Specifically, we adopt a checkerboard as our calibration scene and capture data from it using the FARO laser scanner and Polar device. The relative pose between the FARO laser scanner and the LiDAR sensor in our Polar device is achieved by point cloud registration, and that between the FARO laser scanner and the fisheye camera is obtained by establishing color correspondences between the point cloud captured by the FARO sensor and the fisheye image. By connecting the transformation from FARO to LiDAR and from FARO to the fisheye camera, we can finally calculate the extrinsic parameters between the LiDAR and the fisheye camera. We repeat this process twelve times in our implementation to reduce experimental error. It is important to note that FARO was only used to calibrate between the fisheye camera and the LiDAR before we start scanning a garage, and the process only needs to be conducted once.
@@ Appendix B Data Preprocessing and Mesh cleanup
To ensure data privacy, we anonymized vehicle license plate information recorded in the underground garage data collection. Specifically, we used image blurring to hide license plate numbers from identification. This step is essential to protect individual privacy and comply with data protection regulations. To remove the dynamic objects in the collected data, we apply Segment Anything (Kirillov et al., 2023) to the images. For the point cloud data, we estimate whether the points scanned at one LiDAR frame are also observed at other frames, similar to Schauer and Nüchter(2018). For mesh cleaning, we construct a kd-tree from the point cloud of LIV-SLAM and calculate the nearest neighbour distance of each mesh face (k = 1). If the distance exceeds 0.1 m, we remove the face.
@@ Appendix C Details of depth calculation
To accurately calculate depth, we treat each Gaussian as a probability density function. Following (Zwicker et al., 2001a), we get the depth of each Gaussian primitive, $d_{i}$, by calculating the expected value of depth under this probability density as:
@@
where $g_{i}$ is the Gaussian function, $\boldsymbol{x}=[x_{0},x_{1},x_{2}]^{T}$ is the coordinate of the space in the ray space, where $x_{0},x_{1}$ are the pixel coordinates, $x_{2}$ is oriented in the same direction as the pixel ray. As in (Zwicker et al., 2001a), $t=x_{2}/l$ is the depth of $\boldsymbol{x}$ in the camera coordinate system and $l=\sqrt{x_{0}^{2}+x_{1}^{2}+1}$. We denote the center of the Gaussian in ray space as $\boldsymbol{p}=[p_{0},p_{1},p_{2}]$ and the covariance matrix as $\boldsymbol{\boldsymbol{\Sigma}}$, by defining $\boldsymbol{y}=\boldsymbol{x}-\boldsymbol{p}$, the Eqn. 9 can be simplified as:
@@
(11) $g(\boldsymbol{y}+\boldsymbol{p})=e^{-Ay_{2}^{2}+By_{2}+C}=e^{C}\cdot e^{-Ay_{2}^{2}+By_{2}},$
@@
(12) $\begin{split}A&=\frac{1}{2}\left(\boldsymbol{\Sigma}^{-1}\right)_{2,2}\\ B&=-\left(\boldsymbol{\Sigma}^{-1}\right)_{2,0}y_{0}-\left(\boldsymbol{\Sigma}^{-1}\right)_{2,1}y_{1}\\ C&=-\frac{1}{2}\left(\boldsymbol{\Sigma}^{-1}\right)_{0,0}y_{0}^{2}-\frac{1}{2}\left(\boldsymbol{\Sigma}^{-1}\right)_{1,1}y_{1}^{2}-\left(\boldsymbol{\Sigma}^{-1}\right)_{1,0}y_{0}y_{1}\\ .\end{split}$
@@
(16) $R(D)=\begin{cases}\frac{1}{2\beta}D&(D<\beta)\\ 1-\frac{\beta}{2D}&(D\geq\beta)\end{cases},$